# version 1.0

### v6

@@ 9.3.2 Linear Dynamical Predictors
Suppose that $\|B\|+\|C\|\leq\gamma$, and $\|\mathbf{u}_{t}\|\leq U$ for all $t\in[T]$. Then the class $\prod_{h}^{\star}$ for $h=\gamma\log\frac{T\gamma^{3}U}{\varepsilon}$, $\varepsilon$-approximates the class $\prod^{\star}$.
@@
Consider any sequence and any predictor from $\prod^{\star}$, say $\pi_{1}=\pi_{A,B,C,\mathbf{x}_{0}}$. Consider the predictor $\pi_{2}=\pi_{A,B,C,h}\in\prod_{h}^{\star}$ for the same system matrices. Then we have
$\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|$
$=\left\|\sum_{i=1:t-1}CA^{i-1}B\mathbf{u}_{t-i}+CA^{t-1}\mathbf{x}_{0}-\sum_{i=1:h}CA^{i-1}B\mathbf{u}_{t-i}\right\|$
$=\left\|\sum_{i=h}^{T}CA^{i}B\mathbf{u}_{t-i-1}\right\|$
$\leq\sum_{i=h}^{T}\kappa(1-\delta)^{i}\|C\|\|B\|\|\mathbf{u}_{t-i-1}\|$
$\leq\frac{\gamma^{2}\kappa U}{\delta}(1-\delta)^{h}\leq\frac{\gamma^{2}\kappa U}{\delta}e^{-\delta h},$
for some $\frac{\kappa}{\delta}\leq\gamma$. Using the value of $H$, and summing up over all $T$ iterations, we have
$\sum_{t}\|\mathbf{y}_{t}^{\pi_{1}}-\mathbf{y}_{t}^{\pi_{2}}\|$ $\leq\sum_{t}\frac{\gamma^{2}\kappa U}{\delta}e^{-\delta h}=\varepsilon.$
@@ 9.4 Efficient Learning of Linear Predictors
The previous section gives several natural classes of predictors as well as some of the relationships between them. In particular, the class of linear dynamical predictors is optimal for certain scenarios and requires knowledge of only three matrices. However, the parameterization of this class is nonconvex in the matrices $A,B,C$, which presents a computational challenge.
This challenge is circumvented by using the fact that the class $\prod^{L}$ $\varepsilon$-approximates the class $\prod^{\star}$. In addition, the natural parameterization of $\prod^{L}$ is convex and thus can be learned using online convex optimization.
The following algorithm learns the class of linear predictors in the sense that it attains a sublinear regret bound for any sequence of adversarial cost functions, as well as perturbations and observation noise.
Algorithm 6 Learning $\prod^{L}$ by Online Gradient Descent
1: Input: $M^{1}=M^{1}_{1,1:h},M^{1}_{2,1:k}$, convex constraints set $\mathcal{K}\subseteq{\mathbb{R}}^{h\times d_{y}+k\times d_{u}}$
@@
3:  Predict $\hat{\mathbf{y}}_{t}(M^{t})=\sum_{i=1}^{h}M^{t}_{i}\mathbf{u}_{t-i}+\sum_{j=1}^{k}M^{t}_{j}\mathbf{y}_{t-j}$ and observe true $\mathbf{y}_{t}$.
4:  Suffer loss $\ell_{t}(\hat{\mathbf{y}}_{t}(M^{t}),\mathbf{y}_{t})$, and update:
$M^{t+1}=M^{t}-\eta_{t}\nabla\ell_{t}(M^{t})$
$M^{t+1}=\mathop{\Pi}_{\mathcal{K}}(M^{t+1})$
@@
This algorithm is an instance of the Online Gradient Descent algorithm, and Theorem 6.1 directly gives the following bound for it. We denote by $D,G$ the diameter of the set $\mathcal{K}$ as in Algorithm 6, and by $G$ an upper bound on the norm of the gradients of the loss functions $\ell_{t}$.
@@
For choice of step sizes $\eta_{t}=\frac{D}{G\sqrt{t}}$, and $\mathcal{A}$} being Algorithm 6, we have that
$\mathrm{{regret}}({\mathcal{A}})=\sum_{t}\ell(\hat{\mathbf{y}}_{t}^{\mathcal{A}},\mathbf{y}_{t})-\min_{\pi\in\prod^{L}}\ell(\hat{\mathbf{y}}_{t}^{\pi},\mathbf{y}_{t})\leq 2GD\sqrt{T}.$
@@ 9.5 Summary
We began by formulating the problem in terms of partially observable LDS, highlighting the challenges posed by unknown system matrices and noisy observations. We discussed key approaches for online learning in dynamical systems, emphasizing regret minimization as a performance metric. The regret framework allows direct comparisons between learning algorithms by evaluating their cumulative prediction error against the best fixed predictor in hindsight.
A key takeaway from this chapter is that learning in LDS requires balancing expressivity and computational efficiency. Simple linear predictors may be computationally efficient but lack expressivity, whereas more sophisticated dynamical predictors require careful algorithmic design to remain tractable. The regret analysis provided insights into how online learning algorithms adapt to unknown system parameters while ensuring competitive performance.
This discussion sets the stage for the next chapter on Kalman filtering, where we transition from open-loop predictors to closed-loop estimation techniques that integrate feedback from observations to refine state estimates.
@@ 9.6 Bibliographic Remarks
For a comprehensive treatment of online learning in games and online convex optimization, see [CBL06, Haz16].
@@
The class $\prod^{*}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [HSZ17]. That paper showed how to learn partially observable and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
Under probabilistic assumptions, the work of [HMR18] shows that the matrices $A,B,C,D$ of an unknown and partially observed time-invariant linear dynamical system can be learned using first-order optimization methods. Learning without recovery was studied in [GLS+20] even for marginally stable systems. More recently, tensor methods were used in [BLMY23a], and in [BLMY23b] to learn a mixture of linear dynamical systems.
@@ 9.7 Exercises
9.1. Prove that a $\varepsilon$-approximation of prediction classes implies an additional $\varepsilon T$ regret, as formalized in the Lemma 9.2.
9.2. Prove that the class $\prod^{*}$ is optimal for fully observed linear dynamical systems with zero mean noise.
9.3. Show that the policy class $\prod^{\star}_{h}$ can be learned by the policy class $\prod^{L}_{\tilde{h},k,\kappa}$ with $\tilde{h}=0$.
9.4. Give precise regret bounds for learning the class $\prod^{L}_{h,k,\kappa}$ and $\prod^{\star}_{h}$ as a function of the natural class parameters $h,k,\kappa$ and any other relevant parameter.
9.5. Consider the square loss in Algorithm 6. Give an algorithm based on the Online Newton Step algorithm (see [Haz16]) and write down the resulting regret bound.
9.6. Write down the predictor class $\prod^{\star}$ and $\prod^{\star}_{h}$ for time-varying linear dynamical systems. Research the notion of adaptive regret from [Haz16], and reason about the conditions under which the class $\prod^{L}$ approximates $\prod^{*}$ for time-varying linear dynamical systems.
@@
State estimation plays a central role in control and prediction tasks, where an agent seeks to infer latent states from noisy observations. The Kalman filter is a classical and widely used method for solving this problem, providing optimal state estimates under the assumption of linear system dynamics and Gaussian noise.
A key feature of the Kalman filter and a fundamental limitation is that it requires knowledge of the system matrices, including both the state transition matrix and the observation model. These matrices encode how the system evolves over time and how observations relate to the hidden state. Although the Kalman filter is provably optimal when these parameters are known, real-world applications often involve partially known or entirely unknown system dynamics, making explicit identification challenging.
In contrast, online learning approaches, such as those introduced in Chapter 9, provide a model-free open-loop alternative by directly learning optimal linear predictors from the data. Instead of estimating a latent state and system matrices, these methods construct forecasting models using past observations, making them more adaptive to unknown or time-varying systems. However, these approaches do not incorporate feedback, meaning that they operate in an open-loop setting, predicting future states without correcting for new observations.
This chapter moves beyond open-loop prediction to closed-loop state estimation, where past observations continuously refine future predictions. The Kalman filter accomplishes this by recursively updating estimates based on incoming measurements, achieving Bayes-optimal performance under Gaussian noise. This structured feedback mechanism allows for more accurate and robust estimation compared to open-loop methods, particularly in dynamic and uncertain environments.
Beyond classical Kalman filtering, we also introduce a closed-loop online learning method that serves as a convex relaxation of the Kalman filter. This approach retains the Kalman filtering feedback mechanism, while eliminating the need for explicit system identification, making it applicable in adversarial or uncertain environments. By formulating state estimation as an online convex optimization problem, this method provides provable regret guarantees and serves as a bridge between Bayesian filtering and learning-based estimation techniques.
We provide a rigorous introduction to Kalman filtering, including its derivation, optimality properties, and computational aspects. We also discuss its limitations and contrast it with learning-based alternatives, setting the stage for spectral filtering techniques in the next chapter, which further generalize state estimation in unknown dynamical systems.
@@ 10.1 Observable Systems
Observability is a fundamental concept in control theory that determines whether the internal state of a system can be uniquely inferred from its output over time. In the infinite-horizon setting, prediction is meaningful only if observations remain bounded, a condition closely related to system controllability (as discussed in Section 4.3).
A general dynamical system maps an initial state to an infinite sequence of observations. If this mapping is injective, meaning that each initial state corresponds to a unique sequence of observations, then the system is said to be observable. Formally:
@@
A partially observed dynamical system is observable if and only if every sequence of observations uniquely determines the initial hidden state.
A simple example of an unobservable system is given as the following variant of the double integrator from section 2.1.2. Consider the double integrator dynamical system with $\Delta=1$ and partial observability where the observation matrix is given as follows
$A=\begin{bmatrix}1\ 1\\ 0\ 1\end{bmatrix}\ ,\ C=[0\ 1].$
@@
For linear systems, observability can be verified using a rank condition on the Kalman observability matrix, given by:
$\hat{K}(A,C)=\begin{bmatrix}C\\ CA\\ \dots\\ CA^{d_{\mathbf{x}}-1}\end{bmatrix}$.
where $A$ and $C$ are the system matrices that define the state transition and observation model, respectively. Notice that in our example, the Kalman observability matrix is given by $\hat{K}(A,C)=\begin{bmatrix}0\ 1\\ 0\ 1\end{bmatrix}$, which is degenerate. This is not by chance; the system is observable if and only if the observability matrix has full rank. We ask the reader to provide a proof of the following result in Exercise 10.6.
@@
A partially-observed dynamical system specified by system parameters $(A,B,C)$ is observable if and only if $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))=d_{\mathbf{x}}$.
10.2 Optimality of the Kalman Filter
10.2.1 The Power of Linear Predictors
One of the important conclusions from this chapter is that the policy class of linear predictors is optimal with respect to closed-loop policies. We prove this by proving two facts:
  1. 1.
    The class of linear policies $\varepsilon$-approximates the class of Kalman filters, in the sense of Definition 9.1.
  2. 2.
    The class of Kalman filters is optimal for closed-loop policies.
The conclusion from these two facts is that algorithm 6 can be used to learn a policy that competes with the best closed-loop optimal predictor in hindsight. If the signal comes from a linear dynamical system subject to Gaussian noise, this means competing with the optimal predictor, which is a Kalman filter. We prove this formally at the end of this section.
10.2.2 Closed-Loop Optimal Predictors
Open-loop predictors, such as those introduced in the previous chapter, leverage system inputs and initial conditions to make predictions. However, they do not incorporate feedback from observations, making them susceptible to modeling errors and disturbances.
Closed-loop prediction rules address this limitation by continuously adjusting estimates based on new observations. The most well-known example of such a predictor is the Kalman filter, which produces the best linear predictor in a least-squares sense under the assumption that process and observation noise are Gaussian with known covariances.
Formally, let $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ denote the process and observation noise, assumed to be zero-mean and i.i.d., with covariance matrices:
$\Sigma_{x}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[\mathbf{w}_{t}\mathbf{w}_{t}^{\top}\right],\quad\Sigma_{y}=\mathbb{E}\left[\mathbf{v}_{t}\mathbf{v}_{t}^{\top}\right].$
So far, we have focused on predicting observations, that is, $\mathbf{y}_{t}$. However, the Kalman filter also provides optimal state estimates $\hat{\mathbf{x}}_{t}$ given past observations $\mathbf{y}_{t-1:1}$ and controls $\mathbf{u}_{t-1:1}$. This is motivated by the fact that the least-squares optimal predictor of observations can be expressed as:
$\mathbf{y}_{t}=C\mathbf{x}_{t}.$
In Exercise 10.6, we ask the reader to provide a proof of this statement.
Based upon this observation, the following theorem is the essence of the Kalman filter, which recovers estimates for states as well as linear projection matrices for optimal prediction of the next observation.

### v8

@@ 9.4.1 Convex Relaxation of Stable Systems
Fix $M\geq 1$, and define
$\Pi^{\star}_{\mathrm{stab}}(M)\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\left\{\pi_{\{A,B,C\}}\in\Pi^{\star}:\|A^{t}\|\leq Me^{-t/M}\ \text{for all }t\geq 0,\quad\|B\|+\|C\|\leq M\right\}.$
Assume that $\mathbf{x}_{0}=0$ and that
$\max_{t\in[T]}\|\mathbf{u}_{t}\|\leq M.$
Then $\Pi_{h}^{\star}$ $\varepsilon$-approximates $\Pi^{\star}_{\mathrm{stab}}(M)$ over horizon $T$ provided
$h\geq M\log\!\left(\frac{2M^{5}}{\varepsilon}\right).$
@@
Fix any predictor $\pi_{\{A,B,C\}}\in\Pi^{\star}_{\mathrm{stab}}(M)$, and let $\pi_{\{A,B,C,h\}}\in\Pi_{h}^{\star}$ be its truncated predictor. For each $t\in[T]$,
$\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C\}}}-\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C,h\}}}$ $=\sum_{i=h+1}^{t}CA^{i-1}B\mathbf{u}_{t-i}.$
Hence
$\left\|\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C\}}}-\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C,h\}}}\right\|$ $\leq\sum_{i=h+1}^{t}\|C\|\,\|A^{i-1}\|\,\|B\|\,\|\mathbf{u}_{t-i}\|$
$\leq M^{4}\sum_{i=h+1}^{t}e^{-(i-1)/M}.$
Since $M\geq 1$, we have $1-e^{-1/M}\geq\frac{1}{2M}$, and therefore
$\sum_{i=h+1}^{t}e^{-(i-1)/M}\leq\sum_{i=h+1}^{\infty}e^{-(i-1)/M}\leq\frac{e^{-h/M}}{1-e^{-1/M}}\leq 2Me^{-h/M}.$
Thus
$\left\|\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C\}}}-\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C,h\}}}\right\|\leq 2M^{5}e^{-h/M}.$
Summing over $t=1,\dots,T$ yields
$\sum_{t=1}^{T}\left\|\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C\}}}-\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C,h\}}}\right\|\leq 2TM^{5}e^{-h/M}.$
Choosing
$h\geq M\log\!\left(\frac{2M^{5}}{\varepsilon}\right)$
makes the right-hand side at most $T\varepsilon$, proving the claim. ∎
9.4.2 Exact Representation via Cayley Hamilton
In contrast to the previous approximation result, which is dimension-free but requires stability, the Cayley–Hamilton argument below gives an exact finite-memory representation for arbitrary noiseless finite-dimensional systems, at the price of dependence on the hidden-state dimension $d_{x}$.
For a given matrix $A$ that describes the state evolution of a linear dynamical system with hidden-state dimension $d_{x}$, let
$p_{A}(\lambda)=\lambda^{d_{x}}+a_{1}\lambda^{d_{x}-1}+\cdots+a_{d_{x}}$
be the characteristic polynomial of $A$, and define $a_{0}=1$. The following theorem is the key structural observation of this chapter.
Theorem 9.5.
Consider the noiseless linear dynamical system
$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t},$
$\mathbf{y}_{t}$ $=C\mathbf{x}_{t},$
with characteristic polynomial $p_{A}(\lambda)$. Then there exist matrices
$M_{1,i}\in{\mathbb{R}}^{d_{y}\times d_{y}},\qquad M_{2,j}\in{\mathbb{R}}^{d_{y}\times d_{u}},\qquad i,j=1,\ldots,d_{x},$
such that for every $t>d_{x}$,
$\mathbf{y}_{t}=\sum_{i=1}^{d_{x}}M_{1,i}\mathbf{y}_{t-i}+\sum_{j=1}^{d_{x}}M_{2,j}\mathbf{u}_{t-j}.$
In particular, one may take
$M_{1,i}=-a_{i}I_{d_{y}},\qquad M_{2,j}=C\left(\sum_{i=0}^{j-1}a_{i}A^{j-1-i}\right)B.$
Proof.
Write $d=d_{x}$. By unrolling the state dynamics for $d$ steps, we obtain
$\mathbf{x}_{t}=A^{d}\mathbf{x}_{t-d}+\sum_{j=1}^{d}A^{j-1}B\mathbf{u}_{t-j}.$ (9.1)
Similarly, for each $i=1,\ldots,d$,
$\mathbf{x}_{t-i}=A^{d-i}\mathbf{x}_{t-d}+\sum_{j=i+1}^{d}A^{j-i-1}B\mathbf{u}_{t-j},$
so
$A^{d-i}\mathbf{x}_{t-d}=\mathbf{x}_{t-i}-\sum_{j=i+1}^{d}A^{j-i-1}B\mathbf{u}_{t-j}.$
By the Cayley–Hamilton theorem, $A^{d}+a_{1}A^{d-1}+\cdots+a_{d}I=0$, hence $A^{d}=-\sum_{i=1}^{d}a_{i}A^{d-i}$. Substituting this into (9.1), and then using the previous identity, gives
$\mathbf{x}_{t}$ $=-\sum_{i=1}^{d}a_{i}A^{d-i}\mathbf{x}_{t-d}+\sum_{j=1}^{d}A^{j-1}B\mathbf{u}_{t-j}$
$=-\sum_{i=1}^{d}a_{i}\mathbf{x}_{t-i}+\sum_{i=1}^{d}a_{i}\sum_{j=i+1}^{d}A^{j-i-1}B\mathbf{u}_{t-j}+\sum_{j=1}^{d}A^{j-1}B\mathbf{u}_{t-j}$
$=-\sum_{i=1}^{d}a_{i}\mathbf{x}_{t-i}+\sum_{j=1}^{d}\left(A^{j-1}+\sum_{i=1}^{j-1}a_{i}A^{j-1-i}\right)B\mathbf{u}_{t-j}$
$=-\sum_{i=1}^{d}a_{i}\mathbf{x}_{t-i}+\sum_{j=1}^{d}\left(\sum_{i=0}^{j-1}a_{i}A^{j-1-i}\right)B\mathbf{u}_{t-j}.$
Multiplying both sides by $C$ yields
$\mathbf{y}_{t}=-\sum_{i=1}^{d}a_{i}\mathbf{y}_{t-i}+\sum_{j=1}^{d}C\left(\sum_{i=0}^{j-1}a_{i}A^{j-1-i}\right)B\mathbf{u}_{t-j},$
which is exactly the claimed recurrence. ∎
The theorem shows that every noiseless $d_{x}$-dimensional linear dynamical system induces an exact linear predictor of memory at most $d_{x}$. Thus, if an upper bound on the hidden-state dimension is known, then it suffices to compete with a class of linear predictors of that memory length.
Corollary 9.6.
Let $\pi_{\{A,B,C,\mathbf{x}_{0}\}}\in\Pi^{\star}$ be any noiseless linear dynamical predictor with hidden-state dimension at most $d_{x}$, and let $\{\mathbf{y}_{t}\}_{t=1}^{T}$ be the output sequence it generates under an arbitrary input sequence $\{\mathbf{u}_{t}\}_{t=1}^{T}$. Then there exists $\kappa<\infty$ and a predictor $\pi\in\Pi^{L}_{d_{x},d_{x},\kappa}$ such that
$\hat{\mathbf{y}}_{t}^{\pi}=\hat{\mathbf{y}}_{t}^{\pi_{\{A,B,C,\mathbf{x}_{0}\}}}=\mathbf{y}_{t}\qquad\text{for all }t>d_{x}.$
Consequently, along every realizable trajectory of a noiseless $d_{x}$-dimensional linear dynamical system, the class $\Pi^{L}_{d_{x},d_{x},\kappa}$ exactly represents the induced linear dynamical predictor after a warm-up of $d_{x}$ rounds.
Proof.
Theorem 9.5 explicitly constructs matrices $M_{1,1:d_{x}}$ and $M_{2,1:d_{x}}$ that realize the predictor as a linear predictor of memory $d_{x}$ for all times $t>d_{x}$. Choosing $\kappa$ larger than the corresponding sum of matrix norms yields the claim. ∎
This exact finite-memory representation is one of the main reasons linear predictors are useful. It says that, in the noiseless setting, learning a linear predictor is not merely a heuristic surrogate for learning an unknown linear dynamical system: after an initial transient of length at most $d_{x}$, it is expressive enough to represent the predictor exactly.
In the noisy setting, the same argument gives the same linear recurrence plus an additive disturbance term that depends on recent process and observation noises. Thus, linear predictors remain a natural benchmark even in the stochastic case, though they are no longer exact.
@@ 9.5 Efficient Learning of Linear Predictors
The previous section gives several natural classes of predictors as well as some of the relationships between them. In particular, the class of linear dynamical predictors is natural but nonconvex in the matrices $A,B,C$. The Cayley–Hamilton theorem circumvents this challenge: every noiseless $d_{x}$-dimensional LDS induces a predictor in the convex class $\Pi^{L}_{d_{x},d_{x},\kappa}$ for an appropriate $\kappa$.
The following algorithm learns the class of linear predictors in the sense that it attains a sublinear regret bound for any sequence of convex loss functions.
Algorithm 6 Learning $\Pi^{L}$ by Online Gradient Descent
1: Input: initial predictor matrices $M^{1}=(M^{1}_{1,1:h},M^{1}_{2,1:k})\in\mathcal{K}$, convex constraint set $\mathcal{K}$
@@
3:  Predict
$\hat{\mathbf{y}}_{t}(M^{t})=\sum_{i=1}^{h}M^{t}_{1,i}\mathbf{y}_{t-i}+\sum_{j=1}^{k}M^{t}_{2,j}\mathbf{u}_{t-j},$
and observe true $\mathbf{y}_{t}$.
4:  Suffer loss $\ell_{t}(\hat{\mathbf{y}}_{t}(M^{t}),\mathbf{y}_{t})$ and update
$M^{t+\frac{1}{2}}=M^{t}-\eta_{t}\nabla\ell_{t}(M^{t})$
$M^{t+1}=\mathop{\Pi}_{\mathcal{K}}(M^{t+\frac{1}{2}})$
@@
Here $\ell_{t}(M)$ is shorthand for the loss incurred by predictor $M$ at time $t$, namely $\ell_{t}(\hat{\mathbf{y}}_{t}(M),\mathbf{y}_{t})$.
This algorithm is an instance of the Online Gradient Descent algorithm, and Theorem 6.1 directly gives the following bound for it. We denote by $D$ the diameter of the set $\mathcal{K}$ and by $G$ an upper bound on the norm of the gradients of the loss functions $\ell_{t}(M)$ over $\mathcal{K}$.
@@
For the choice of step sizes $\eta_{t}=\frac{D}{G\sqrt{t}}$, and $\mathcal{A}$} being Algorithm 6, we have
$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi^{L}_{h,k,\kappa})\leq 2GD\sqrt{T}.$
Combining Corollary 9.6 with Corollary 9.7, we conclude that if the observations are generated by an unknown noiseless linear dynamical system of hidden-state dimension at most $d_{x}$, then learning $\Pi^{L}_{d_{x},d_{x},\kappa}$ suffices to obtain sublinear regret with respect to the best linear dynamical predictor in hindsight, up to the lower-order cost of the first $d_{x}$ initialization rounds.
@@ 9.6 Summary
We began by formulating the problem in terms of partially observable LDS, highlighting the challenges posed by unknown system matrices and noisy observations. We then introduced several natural classes of predictors and explained how to compare them through regret minimization.
The key technical observation of the chapter is that every noiseless finite-dimensional linear dynamical system induces an exact finite-memory linear predictor. This follows from the Cayley–Hamilton theorem, and it explains why simple linear predictors are far more expressive than they might initially appear. Rather than identifying the unknown system matrices explicitly, we can compete directly with the induced input-output predictor.
Finally, we showed that linear predictors can be learned efficiently by online convex optimization, using Online Gradient Descent to obtain sublinear regret.
@@ 9.7 Bibliographic Remarks
Learning linear predictors with finite history is a classical consequence of the Cayley–Hamilton theorem and realization theory; see, e.g., [KAI80, CHE99].
For a comprehensive treatment of online learning in games and online convex optimization, see [CL06, HAZ16].
@@
The class $\Pi^{\star}$, which is the class of optimal open-loop predictors for systems with full observation, was studied in [HSZ17]. That paper showed how to learn partially observable and even marginally stable dynamical systems whose spectral radius approaches one, as long as their transition matrix is symmetric.
Under probabilistic assumptions, the work of [HMR18] shows that the matrices $A,B,C,D$ of an unknown and partially observed time-invariant linear dynamical system can be learned using first-order optimization methods. Learning without recovery was studied in [GLS+20] even for marginally stable systems. More recently, tensor methods were used in [BLM+23a], and in [BLM+23b] to learn a mixture of linear dynamical systems.
The idea that a finite-dimensional state-space model induces a finite-order input-output recurrence is classical in linear systems theory. In the context of this chapter, the Cayley–Hamilton theorem provides a particularly clean explanation for why linear predictors can exactly represent noiseless finite-dimensional LDS.
@@ 9.8 Exercises
9.1. Prove that Lemma 9.2 holds.
9.2. Assume the system is fully observed, so that $C=I$, and that the process noise has zero conditional mean given the past. Show that the open-loop predictor obtained by unrolling the dynamics minimizes the expected squared prediction error among all predictors measurable with respect to the past inputs.
9.3. Show that every truncated open-loop predictor of the form
$\hat{\mathbf{y}}_{t}=\sum_{i=1}^{h}H_{i}\mathbf{u}_{t-i}$
is a special case of a linear predictor in the class $\Pi^{L}_{0,h,\kappa}$ for an appropriate choice of $\kappa$.
9.4. Assume that $\|\mathbf{u}_{t}\|\leq U$, $\|\mathbf{y}_{t}\|\leq Y$ for all $t$, and that the parameter space defining $\Pi^{L}_{h,k,\kappa}$ is endowed with a norm under which its diameter is at most $D$. Derive an explicit upper bound on $\mathrm{{regret}}_{T}({\mathcal{A}},\Pi^{L}_{h,k,\kappa})$ for Online Gradient Descent in terms of $h,k,\kappa,U,Y,D$, and $T$.
9.5. Assume the square loss, and suppose moreover that predictions, outputs, and the parameter domain are uniformly bounded. Give an algorithm based on Online Newton Step for learning $\Pi^{L}_{h,k,\kappa}$, and derive the corresponding regret bound.
@@ 9.3. Show that every truncated open-loop predictor of the form
In the previous chapter we defined what it means to learn in dynamical systems, and established relationship between prediction classes. Most notably, we showed that despite nonconvexity of the class of linear dynamical predictors, they can be learned exactly using the convex relaxation of linear predictors.
In this chapter we show another approach to learn the class of linear dynamical predictors. This approach is more limited in terms of the generality where it applies, it is based on generative data and statistical assumptions about the noise. However, it allows for more than just prediction: Kalman filtering allows state estimation which is very useful by its own right in various applications.
@@ 10.1 Observable Systems
Observability asks whether the hidden state of a partially observed dynamical system can be recovered from its observations. This is the basic structural condition that makes state estimation meaningful: if two different hidden states can produce the same observation history, then no estimator—Kalman filter or otherwise—can determine the state uniquely.
This notion is dual to controllability, but conceptually different. Controllability asks whether one can drive the state by choosing controls; observability asks whether one can infer the state from what is observed. For linear systems, observability admits a clean rank characterization through the Kalman observability matrix.
@@
Consider the noiseless partially observed linear system
$\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t},\qquad\mathbf{y}_{t}=C\mathbf{x}_{t},$
with known input sequence $\{\mathbf{u}_{t}\}_{t\geq 0}$. We say that the pair $(A,C)$ is observable if for every two initial states $\mathbf{x}_{0},\mathbf{x}_{0}^{\prime}$, equality of the resulting observation sequences under the same inputs implies $\mathbf{x}_{0}=\mathbf{x}_{0}^{\prime}$. Equivalently, the observation history uniquely determines the initial state.
A simple example of an unobservable system is the following variant of the double integrator from Section 2.1.2. Consider the double integrator dynamical system with $\Delta=1$ and partial observability where the observation matrix is given by
$A=\begin{bmatrix}1&1\\ 0&1\end{bmatrix},\qquad C=\begin{bmatrix}0&1\end{bmatrix}$.
@@
For linear systems, observability can be verified using a rank condition on the Kalman observability matrix, given by
$\hat{K}(A,C)\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\begin{bmatrix}C\\ CA\\ \dots\\ CA^{d_{\mathbf{x}}-1}\end{bmatrix}$.
where $A$ and $C$ are the system matrices that define the state transition and observation model, respectively. Notice that in our example, the Kalman observability matrix is given by
$\hat{K}(A,C)=\begin{bmatrix}0&1\\ 0&1\end{bmatrix}$,
which is degenerate. This is not by chance: the system is observable if and only if the observability matrix has full rank. We ask the reader to provide a proof of the following result in Exercise 10.6.
@@
A partially-observed dynamical system with system matrices $(A,C)$ is observable if and only if $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))=d_{\mathbf{x}}$.
10.2 The Kalman Filter
Chapter 9 focused on open-loop predictors, which map past observations and controls to a forecast of the next output. Such predictors do not maintain or update an explicit estimate of the hidden state. The Kalman filter is the canonical closed-loop predictor for linear systems with zero-mean process and observation noise. It recursively combines the system dynamics with the latest measurement, producing the best linear least-squares estimate of the hidden state when the system matrices and noise covariances are known.
Formally, let $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ denote the process and observation noise, assumed to be zero-mean and i.i.d., with covariance matrices
$\Sigma_{x}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}[\mathbf{w}_{t}\mathbf{w}_{t}^{\top}],\qquad\Sigma_{y}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}[\mathbf{v}_{t}\mathbf{v}_{t}^{\top}].$
Although our ultimate goal is to predict observations, it is convenient to first estimate the hidden state. Once we have a least-squares estimate $\hat{\mathbf{x}}_{t}$ of $\mathbf{x}_{t}$ given the past, the corresponding least-squares prediction of the observation is $\hat{\mathbf{y}}_{t}=C\hat{\mathbf{x}}_{t}$; see Exercise 10.6.
The following theorem is the essence of the Kalman filter. It constructs a recursive state estimator that is optimal among linear estimators based on past observations and controls.